### v3

# The NP-hardness of the Gromov-Wasserstein distance

Abstract
This note addresses the property frequently mentioned in the literature that the Gromov-Wasserstein (GW) distance is NP-hard. We provide the details on the non-convex nature of the GW optimization problem that imply NP-hardness of the GW distance between finite spaces for any instance of an input data. We further illustrate the non-convexity of the problem with several explicit examples.
The Gromov-Wasserstein (GW) distance defined in [21, 22] accepts a pair of metric-measure spaces and outputs a distance between them111up to measure-preserving isometries, which makes GW a metric on the equivalence classes of metric-measure spaces, see Theorem 5.1(a) of [22].. This metric property makes GW distance highly applicable for comparison of objects, in particular objects that lie in different spaces, - a desirable task in scientific and machine learning applications. Scientific applications include aligning density maps in cryo-EM [27], comparing molecular structures in chemistry [25], inferring cell-cell communications [6] and aligning data from different experiments [14] in biology, and classifying biomedical time series [19], to name a few. Machine learning applications pertain to generative modeling [5], reinforcement learning [16], and graph partitioning and clustering [11, 31, 33], among others.
In order to perform GW based object comparison, one needs to compute a GW distance, a constrained minimization problem frequently mentioned as being NP-hard due to its non-convexity (see, e.g., discussion in Section 7 of [22], Section 1.1 of [25], Section 1 of [29], and/or Section 3 of [8]). To solve this problem in practice, two main approaches are currently taken in the literature, none of which guarantees a global minimum. The first approach is to employ the conditional gradient method (also known as Frank-Wolfe), which converges to a stationary point of the objective function that may not be a global minimum in the non-convex case (see Sections 2.1 and 2.2.2 of [4]). Conditional gradient for GW distance is implemented in the Python toolbox [17] and used for variants of GW problem in [7, 2]. The second approach proposed in [25] is to regularize the objective by adding an entropy term (see equation (7) of [25] and discussion in Section 10.6.4 in [24]). Such regularization does not make the problem convex222and hence no global minima are guaranteed when solving the problem via projected gradient descent proposed in [25] - see Remark 3 of [25], but allows for efficient implementations (Python toolboxes [17] and [12]) making the regularization technique widely used for various GW related problems [34, 9, 20, 28].
This work provides the details on the structure of the GW optimization problem for finite spaces333the case most commonly encountered in real life applications that permit establishing the non-convex quadratic nature and imply NP-hardness of the GW distance. We note that the value of a GW distance is analytically determined in several special cases of metric-measure spaces with infinitely many points: see Theorem 4.1 [13] for GW distance between Gaussians when restricting the problem to Gaussian plans, and Theorem 1 of [1] for GW distance between spheres. For finite spaces, [23] proposes a polynomial time algorithm to compute an ultrametric variant of the GW distance with exponent $p=\infty$ (see Section 5.1 of the reference). Nevertheless, as the current note highlights, computing the $p$-GW distance, $p\in[0,\infty)$, between two datasets with finitely many points remains an NP-hard problem for any instance of an input data.
The remainder of this note is organized as follows: Definition 1 defines a GW distance as an optimization program, followed by Example 3 that illustrates the structure of the program and its non-convexity on a pair of small-size exemplar metric spaces. Theorem 4 (main result) shows that GW distance between any two metric-measure spaces with finitely many points is NP-hard to compute. The result of the theorem is applied in Observation 5 and illustrated on the real life data in Example 6. The datasets and codes to reproduce Example 6 are available at https://github.com/kravtsova2/GW_NPhard.
Definition 1 (GW distance (finite spaces case), Definition 5.7 of [22]).
Given two finite metric-measure spaces $\mathcal{X}:=(X,d_{X},\mu_{X})$ and $\mathcal{Y}:=(Y,d_{Y},\mu_{Y})$ with points $X=\{x_{1},\ldots,x_{m}\}$ and $\{y_{1},\ldots,y_{n}\}$ respectively, define $GW_{p}^{p}(\mathcal{X},\mathcal{Y})$, $p\in[1,\infty)$, by the optimal value of the optimization problem444see also equation (4) of [25] and/or equations (1) and (2) of [29] that use the same notation as we use here.
$\min_{\mu\in\mathcal{C}(\mu_{X},\mu_{Y})}\sum_{i,j,k,l}(\Gamma_{p})_{ijkl}\cdot\mu_{ij}\mu_{kl}$ (1)
where
  • The objective coefficients are stored in a 4-way tensor $\Gamma_{p}$ with elements $(\Gamma_{p})_{ijkl}=\left|d_{X}(x_{i},x_{k})-d_{Y}(y_{j},y_{l})\right|^{p}$, $i,k\in\{1,\ldots,m\}$, $j,l\in\{1,\ldots,n\}$.
  • The constraint set $\mathcal{C}(\mu_{X},\mu_{Y})$ denotes the set of couplings between $\mu_{X}$ and $\mu_{Y}$, i.e. Borel probability measures on $X\times Y$ with marginals $\mu_{X}$ and $\mu_{Y}$, respectively.
  • The optimization variables $\mu_{i^{\prime}j^{\prime}},i^{\prime}\in\{1,\ldots,m\},j^{\prime}\in\{1,\ldots,n\}$ stored in a vector representing a probability measure $\mu$ satisfying the coupling constraints above.
The p-GW distance between spaces $\mathcal{X}$ and $\mathcal{Y}$ is then given by $\frac{1}{2}\left(GW_{p}^{p}(\mathcal{X},\mathcal{Y})\right)^{1/p}$ [^5] Note that neither taking the $p$th root nor subsequent scaling by $\frac{1}{2}$ affects the nature of the optimization problem; hence, we will be focusing on the object $GW_{p}^{p}(\mathcal{X},\mathcal{Y})$ in what follows..
Observation 2.
Observe that (see Example 3 for a concrete illustration of these quantities)
  1. 1.
    The optimization variable $\mu$ can be written as a long vector (of length $m\cdot n$) with entries $\mu_{i^{\prime}j^{\prime}}$.
  2. 2.
    The tensor $\Gamma_{p}$ can be written in a matrix form by writing the $(i,k)$th block of $\underset{m\cdot n\times m\cdot n}{\Gamma_{p}}$ to be the $n\times n$ matrix with entries $\left|d_{X}(x_{i},x_{k})-d_{Y}(y_{j},y_{l})\right|^{p}$. Note that this matrix is symmetric due to symmetry of the distances $d_{X}$ and $d_{Y}$ [^6] The symmetry is not guaranteed if the distances on the ground spaces are replaced by more general quantities such as asymmetric weight functions on networks [10]..
  3. 3.
    The constraints $\mu\in\mathcal{C}(\mu_{X},\mu_{Y})$ can be written in a matrix form $A\mu=\begin{pmatrix}\mu_{X}\\ \mu_{Y}\end{pmatrix}$ for a matrix $A$ responsible for summing appropriate entries of the joint probability vector $\mu$.
This allows to write the $GW_{p}^{p}(\mathcal{X},\mathcal{Y})$ program (1) as a linearly constrained quadratic optimization program
[…]
$A\mu=b$
The following Example 3 illustrates the structure of the program in (2):
Example 3.
(Spaces are from Example 5.1 of [22])). Consider the space $\mathcal{X}:=(X,d,\mu_{X})$ with two points $X:=\{x_{1},x_{2}\}$ with distance $d$ giving $(x_{1},x_{2})=d(x_{2},x_{1})=1$ and $d(x_{1},x_{1})=d(x_{2},x_{2})=0$ (this can be conveniently represented by a distance matrix $(d_{ik})=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$) and measure $\mu_{X}:=(\frac{1}{2},\frac{1}{2})$. A pictorial representation of such space is provided in Figure 1A, top panel. Similarly, consider the space $\mathcal{Y}:=(Y,h,\mu_{Y})$ with points $Y:=(y_{1},y_{2})$, distance $(h_{jl})=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and measure $\mu_{Y}:=(\frac{1}{4},\frac{3}{4})$.
Note: The space $\mathcal{X}=\Delta_{2}$ is a member of the family of spaces defined in Example 5.2 of [22] as $(\Delta_{n})_{n\in\mathbb{N}}$, where each member $\Delta_{n}$ contains $n$ points and is equipped with the distance $(d_{n})_{ij}=1-\delta_{i,j}$ and the uniform measure $\mu_{n}=(\frac{1}{n},\cdots,\frac{1}{n})$ (the second space $\mathcal{Y}$ above differs in having non-uniform measure). The setting described here is generalized in Example 6(a) where $\Delta_{2}$ is compared with $\Delta_{n}$ for $n=2,\ldots,50$.
The GW distance seeks to find the joint probability measure with marginals $\mu_{X}$ and $\mu_{Y}$, i.e. the vector $\mu:=\begin{pmatrix}\mu_{11}\\ \mu_{12}\\ \mu_{21}\\ \mu_{22}\end{pmatrix}$ satisfying the constraints
$\underbrace{\begin{pmatrix}1&1&0&0\\ 0&0&1&1\\ 1&0&1&0\\ 0&1&0&1\end{pmatrix}}_{A}\underbrace{\begin{pmatrix}\mu_{11}\\ \mu_{12}\\ \mu_{21}\\ \mu_{22}\end{pmatrix}}_{\mu}=\underbrace{\begin{pmatrix}1/2\\ 1/2\\ 1/4\\ 3/4\end{pmatrix}}_{b}$
which minimizes the objective
$\mu^{\prime}\Gamma\mu=\sum_{i,j,k,l}\Gamma_{ijkl}\mu_{ik}\mu_{jl}$
where $\Gamma_{ijkl}=|d_{ik}-h_{jl}|^{p}$. The objective here can be written (in this specific case, it is the same for all $p\in[1,\infty)$) as
$\underbrace{\begin{pmatrix}\mu_{11}&\mu_{12}&\mu_{21}&\mu_{22}\end{pmatrix}}_{\mu^{\prime}}\underbrace{\begin{pmatrix}|d_{11}-h_{11}|&|d_{11}-h_{12}|&|d_{12}-h_{11}|&|d_{12}-h_{12}|\\ |d_{11}-h_{21}|&|d_{11}-h_{22}|&|d_{12}-h_{21}|&|d_{12}-h_{22}|\\ |d_{21}-h_{11}|&|d_{21}-h_{12}|&|d_{22}-h_{11}|&|d_{22}-h_{12}|\\ |d_{21}-h_{21}|&|d_{21}-h_{22}|&|d_{22}-h_{21}|&|d_{22}-h_{22}|\end{pmatrix}}_{\Gamma}\underbrace{\begin{pmatrix}\mu_{11}\\ \mu_{12}\\ \mu_{21}\\ \mu_{22}\end{pmatrix}}_{\mu}$ (3)
The GW distance between $\mathcal{X}$ and $\mathcal{Y}$ is $\frac{1}{2}\left(GW_{p}^{p}(\mathcal{X},\mathcal{Y})\right)^{1/p}$, where $GW_{p}^{p}(\mathcal{X},\mathcal{Y})$ is the optimal value of the quadratic program
$\min_{\mu\geq 0}\mu^{\prime}\Gamma\mu$
$A\mu=b$
For the spaces $\mathcal{X}$ and $\mathcal{Y}$ defined above, the matrix $\Gamma$ reads
$\Gamma=\begin{pmatrix}0&1&1&0\\ 1&0&0&1\\ 1&0&0&1\\ 0&1&1&0\end{pmatrix}$
Observe that the matrix $\Gamma$ has a negative eigenvalue $\lambda=-2$, and hence it is not positive semidefinite.
Theorem 4 below shows that the presence of negative eigenvalues in the GW objective matrix is not specific to Example 3, but rather holds in general when posing GW problem between any pair of metric-measure spaces with finitely many points. The result states that the matrix in the objective function of GW distance from Definition 1 cannot be positive semidefinite, regardless of an input data. Hence, the resulting quadratic program is not convex and thus is NP-hard from computational perspective ([30]; for a recent review on the literature on computational complexity results for quadratic programming, see [15]).
Theorem 4 (GW distance between finite spaces is NP-hard).
The GW distance program in Definition 1 is non-convex quadratic (and hence is NP-hard).
Proof.
Observe that (taking $p=1$ for simplicity), for any general pair of finite metric measures spaces with at least two points and distance matrices $(d_{ik})$ and $(h_{jl})$, a matrix $\underset{m\cdot n\times m\cdot n}{\Gamma}$ as in the objective (2) will always have a submatrix $\Gamma_{[1:2,1:2]}=\begin{pmatrix}|d_{11}-h_{11}|&|d_{11}-h_{12}|\\ |d_{11}-h_{21}|&|d_{11}-h_{22}|\end{pmatrix}$ in the left corner, i.e. the matrix $\Gamma$ is given by
$\Gamma=\begin{pmatrix}|d_{11}-h_{11}|&|d_{11}-h_{12}|&\cdots\\ |d_{11}-h_{21}|&|d_{11}-h_{22}|&\cdots\\ \vdots&\vdots&\ddots\end{pmatrix}$
Since the $(d_{ij})$ and $(h_{ij})$ are the distances, it holds that $d_{11}=h_{11}=h_{22}=0$, while $h_{12}=h_{21}>0$. Hence, the determinant
$\text{det}(\Gamma_{[1:2,1:2]})=0-h_{12}h_{21}<0$
i.e. $\Gamma$ has a negative principal minor (the determinant of a submatrix taken from $\Gamma$ according to index subsets $I\subseteq[m\cdot n]$ and $J\subseteq[m\cdot n]$ with $I=J$). Note that the conclusion would not change if $p>1$, in which case one would still have $\text{det}(\Gamma_{p_{[1:2,1:2]}})=0-h_{12}^{p}h_{21}^{p}<0$.
By Sylvester’s criterion stating that all principal minors of a real symmetric positive semidefinite matrix must be non-negative (see, e.g., [26]), the matrix $\Gamma$ or, more generally, $\Gamma_{p}$, - cannot be positive semidefinite. By the discussion preceding the theorem, a quadratic program with such objective cannot be convex and hence is NP-hard. ∎
Theorem 4 indicates that every instance of the GW distance optimization program is NP-hard, which supports the ongoing efforts on its computation (see, e.g., [23] for a recent review). At the same time, GW distance can itself be viewed as a relaxation of the Gromov-Hausdorff distance problem (see Section 4.1 of [22]) that generalizes a Quadratic Assignment Problem (QAP) (see Remark 4.6 of [22] for the connection between Gromov-Hausdorff and QAP). Below we discuss the interpretation of GW as a relaxation of QAP in the form that QAP was introduced in the economic literature, and use Theorem 4 to argue that such relaxation is not convex.
Observation 5 (Relation to the Quadratic Assignment Problem).
The structure of the problem (2) is related to the (NP-hard) QAP introduced by T. Koopmans and M. Beckmann in [18]. The program describes an optimal assignment of $m$ industrial plants to $m$ locations such that the plants with high flows of commodities between them are placed close to each other (see Section 6 of [18] for the description). The flows between pairs of plants are represented by a square matrix $(f_{ik})$, and the distances between target locations are stored in a square matrix $(h_{jl})$. The problem is to find a $m\times m$ permutation matrix $X_{i^{\prime}j^{\prime}}$ with entries $X_{i^{\prime}j^{\prime}}=1$ if the $i^{\prime}$th plant is placed into the $j^{\prime}$th location, and zero otherwise, which describes an optimal assignment of plants to locations. This problem reads
$\min_{X}\sum_{ijkl}f_{ik}h_{jl}X_{ij}X_{kl}$
$\sum_{j^{\prime}}X_{i^{\prime}j^{\prime}}=1\text{ }\forall i^{\prime}=1,\ldots,n$
$\sum_{i^{\prime}}X_{i^{\prime}j^{\prime}}=1\text{ }\forall j^{\prime}=1,\ldots,n$
$X_{i^{\prime}j^{\prime}}\in\{0,1\}$
Writing $X$ as a long vector with $m^{2}$ entries $X_{i^{\prime}j^{\prime}}$ allows to represent the QAP in the form
$\min_{x\in\{0,1\}}x^{\prime}\Gamma x$ (4)
$Ax=b$
with $\Gamma_{ijkl}=f_{ik}h_{jl}$ (see [3] for this formulation of QAP).
To see the relation between QAP and GW, consider the objective of the GW distance problem in Definition 1 between metric-measure spaces with ground distances $(d_{ij})$ and $(h_{ij})$, under the choice of exponent $p=2$, which reads
$\mu^{\prime}\Gamma_{p}\mu=\sum_{ijkl}(d_{ik}-h_{jl})^{2}\cdot\mu_{ij}\mu_{kl}$
$=\underbrace{\sum_{ijkl}d_{ik}^{2}\cdot\mu_{ij}\mu_{kl}+\sum_{ijkl}h_{jl}^{2}\cdot\mu_{ij}\mu_{kl}}_{\text{depend on $\mu_{X}$ and $\mu_{Y}$ only, and thusare irrelevant for optimization}}-\sum_{ijkl}2d_{ik}h_{jl}\cdot\mu_{ij}\mu_{kl}$
$=\text{constant}+2\sum_{ijkl}\underbrace{(-d_{ik})}_{\text{call }f_{ik}}h_{jl}\cdot\mu_{ij}\mu_{kl}$
The expression above is exactly the QAP objective in (4) if one chooses the flow function $f$ between industrial plants to be the negative distance between them. The optimization variables $\mu$ in GW program are constrained more “softly” to be in the interval $[0,1]$, in contrast to the “hard” QAP constraint for $x$ to take integer values in $\{0,1\}$.
This allows to interpret GW problem as a relaxation of the QAP in a framework conceptually similar to interpreting a Kantorovich problem as a relaxation of the Linear Assignment Problem (LAP) (discussed in Sections 2.2 - 2.3 of [24]) [^7] L. Kantorovich and T. Koopmans shared Nobel Prize in Economics in 1975 for their contributions to optimal resource allocation problems https://www.nobelprize.org/prizes/economic-sciences/1975/press-release/. As can be seen from Theorem 4, the GW relaxation of the QAP is not convex.
We conclude this note with two additional illustrations of the presence of negative eigenvalues in the GW objective matrix, which demonstrates the conclusion of Theorem 4 on synthetic and real world data instances (Example 6 (a) and (b), respectively).
Figure 1: Two additional examples illustrating the presence of negative eigenvalues in $\Gamma_{p}$, as described in Example 6. A. Top panel: a pair of spaces from the family constructed by [22] (Example 5.2). Bottom panel: Fixing the first space with $m=2$ points and varying the number of points $n$ of the second space results in the $\Gamma_{p}$, $p=1$, having increasing number of negative eigenvalues. B. Top panel: a pair of spaces constructed in [19] using the model from [32] ($m=n=500$ points in each space is shown). Bottom panel: Fixing the first space with $m=50$ and varying number of points in the second space results in $\Gamma_{p}$, $p=1$, having increasing number of negative eigenvalues.
Example 6 (Two additional illustrations for negative eigenvalues of $\Gamma_{p}$, Figure 1).
  • (a)
    Spaces from Example 5.2 of [22] This illustration uses the family of spaces $(\Delta_{n})$ from Example 5.2 of [22] and described in Example 3 of this note. Figure 1A, top panel, depicts a sample pair $(X,d_{X},\mu_{X})=\Delta_{2}$ and $(Y,d_{Y},\mu_{Y})=\Delta_{4}$.
    To illustrate a presence of negative eigenvalues in the GW objective matrix $\Gamma_{p}$, we fix the first space to be $\Delta_{2}$, and vary the second space $\Delta_{n}$ by increasing the number of points, $n=2,\ldots,50$. For $p=1$, we report the number of negative eigenvalues in $\Gamma_{p}$, which appears to increase roughly linearly with $n$ (Figure 1A, bottom panel).
  • (b)
    Spaces based on [32] (see [19] for the GW based comparison formulation) Here we consider pairs of parameterized curves in $3D$ whose points are given by $f(t):=(f_{1}(t),f_{2}(t),f_{3}(t))$ and $g(t):=(g_{1}(t),g_{2}(t),g_{3}(t))$, $t\in[0,T]$, respectively. The functions $f$ and $g$ are taken to be the solutions of the system of tree coupled differential equations constructed in [32] (see equation (6) in the reference) corresponding to two different parameter regimes making the equilibrium point (1,1,1) either unstable or stable focus, respectively (Figure 1B, top panel). The problem was considered in [19] as a GW based comparison problem (see Figure 3 and corresponding description in the reference), where one views each trajectory as a metric measure space under the arc length distance and a uniform measure on each space.
    We fix the first space $(X,d_{X},\mu_{X})$ to have $m=50$ sample points from the first trajectory, and vary the second space $(Y,d_{Y},\mu_{Y})$ by sampling $n=10,11,\ldots,50$ points from the second trajectory. For $p=1$, we report the number of negative eigenvalues in $\Gamma_{p}$, which appears to increase roughly linearly with $n$ (Figure 1B, bottom panel).
Acknowledgements and Funding The author thanks Facundo Mémoli for suggesting this problem, providing guidance during the solution process, and supporting the author with NSF DMS #2301359 and NSF CCS #1839356. The author thanks Adriana Dawes for the advice and support. A version of the problem considered in this note was given as an assignment during the topics course Optimal Transport based methods in Data Science (Math8610) taught by Facundo Mémoli at The Ohio State University in the Spring 2023 semester (https://github.com/ndag/OT-DS).